# ActionPiece: Contextually Tokenizing Action Sequences for Generative Recommendation

### v1

@@ Abstract
Generative recommendation (GR) is an emerging paradigm where user actions are tokenized into discrete token patterns and autoregressively generated as predictions. However, existing GR models tokenize each action independently, assigning the same fixed tokens to identical actions across all sequences without considering contextual relationships. This lack of context-awareness can lead to suboptimal performance, as the same action may hold different meanings depending on its surrounding context. To address this issue, we propose ActionPiece to explicitly incorporate context when tokenizing action sequences. In ActionPiece, each action is represented as a set of item features, which serve as the initial tokens. Given the action sequence corpora, we construct the vocabulary by merging feature patterns as new tokens, based on their co-occurrence frequency both within individual sets and across adjacent sets. Considering the unordered nature of feature sets, we further introduce set permutation regularization, which produces multiple segmentations of action sequences with the same semantics. Experiments on public datasets demonstrate that ActionPiece consistently outperforms existing action tokenization methods, improving NDCG@$10$ by $6.00\%$ to $12.82\%$.
@@ 1 Introduction
Generative recommendation (GR) (Geng et al., 2022; Rajput et al., 2023; Zheng et al., 2024a; Zhai et al., 2024) is an emerging paradigm for the sequential recommendation task (Hidasi et al., 2016; Kang & McAuley, 2018). By tokenizing the user actions (typically represented by the interacted items) into discrete tokens, GR models learn to autoregressively generate tokens, which are then parsed into recommended items. These tokens share a compact vocabulary that does not scale with the item pool size, improving model scalability, memory efficiency, and recommendation performance (Rajput et al., 2023; Zhai et al., 2024). The input action sequence is vital in understanding user intentions (Li et al., 2017; Kang & McAuley, 2018), which organizes a user’s historical interactions in chronological order. The same action (e.g., purchasing the same item) may have different meanings in different action sequences. Evidence of taking a certain action can be found in the context, such as whether other items in the sequence share the same brand, color tone, or price range (Zhang et al., 2019; Zhou et al., 2020; Hou et al., 2022, 2023; Yuan et al., 2023).
Figure 1: Illustration of the tokenization process of ActionPiece. Each action is represented as an unordered feature set. This figure presents two possible tokenized sequences. The same action can be tokenized into different tokens depending on the surrounding context. A detailed case study can be found in Section 4.5.
Despite the importance of contextual relations among actions, existing methods tokenize each action independently of its context (summarized in  Table 1). The typical pipeline for tokenizing action sequences involves two steps: (1) Tokenizing each action/item individually into a pattern of tokens; (2) Replacing each action in the input sequence with its corresponding token pattern. In this way, the tokens do not explicitly contain the context. Instead, they solely rely on the autoregressive model’s parameters being well-trained to generalize effectively in understanding the context, which challenges the capabilities of GR models. As a comparison, tokenization in language modeling also originates from context-independent methods, such as word-level tokenization (Sutskever et al., 2014; Bahdanau et al., 2015). A decade of progress has led to most tokenization methods for modern large language models (LLMs) (OpenAI, 2022; Anil et al., 2023; Touvron et al., 2023; Zhao et al., 2023) adopting context-aware approaches, including BPE (Sennrich et al., 2016) and Unigram tokenization (Kudo, 2018), which tokenize the same characters along with their adjacent context into different tokens.
@@
To this end, we propose ActionPiece, which enables the same actions to be tokenized into different tokens based on their surrounding context. (1) Vocabulary construction: We first initialize the vocabulary to include every unique feature as initial tokens. The vocabulary is then constructed by iteratively learning merge rules. Each merge rule specifies that a pair of tokens can be merged into a new token. In each iteration, we enumerate the training corpus to count the co-occurrence of existing tokens. Considering the structural differences between token pairs, e.g., whether they occur within a single set or between two adjacent sets, we assign different weights to different pairs during the counting process. (2) Segmentation: The next step is to segment action sequences into token sequences for GR model training and inference. To fully exploit the unordered nature of the feature set for each action, we introduce set permutation regularization. By randomly permuting the features within each set, we can produce multiple token sequences of a single action sequence that preserve the same semantics. These variations act as natural augmentations for training data and enable inherent ensembling during model inference.
@@ 2 Related Work
Generative recommendation. Conventional sequential recommendation models often relies on large embedding tables to store representations for all items, leading to significant engineering and optimization challenges (Hidasi et al., 2016; Kang & McAuley, 2018). Generative recommendation (Deldjoo et al., 2024; Rajput et al., 2023; Zheng et al., 2024a) addresses these issues by tokenizing each item as tokens from a shared table. By autoregressively generating the next tokens as recommendations, this generative paradigm offers benefits such as memory efficiency (Rajput et al., 2023; Yang et al., 2024; Ding et al., 2024), scalability (Zhai et al., 2024; Liu et al., 2024c), and easier alignment with LLMs (Zheng et al., 2024a; Jin et al., 2024; Tan et al., 2024; Li et al., 2025). Existing research has developed different action tokenization techniques, such as hierarchical clustering (Hua et al., 2023; Si et al., 2024), quantization (Rajput et al., 2023; Wang et al., 2024a; Zhu et al., 2024), or jointly training with recommendation models (Liu et al., 2024a). Other works incorporate additional modalities like collaborative filtering (Petrov & Macdonald, 2023; Wang et al., 2024c, b; Liu et al., 2024c), text (Zheng et al., 2024a; Jin et al., 2024; Hou et al., 2024; Zhang et al., 2025), and vision signals (Liu et al., 2024b). However, current methods tokenize each action independently, ignoring the surrounding context. In this work, we propose the first context-aware action tokenization method, where the same actions are tokenized differently in different action sequences.
@@
3 Method
In this section, we present ActionPiece, a context-aware method for tokenizing action sequences in generative recommendation. First, we formulate the task in Section 3.1. Then, we introduce the proposed tokenizer, covering vocabulary construction and sequence segmentation, in Section 3.2. Finally, we describe the model training and inference process using ActionPiece-tokenized sequences in Section 3.3.
3.1 Problem Formulation
3.2 Contextual Action Sequence Tokenizer
The proposed tokenizer is designed to transform action sequences (represented as sequences of feature sets) into token sequences. In the ActionPiece-tokenized sequences, each token corresponds to a set containing varying numbers of features. For example, a token can represent: (1) a subset of features from one item; (2) a set with a single feature; (3) all features of one item; or (4) features from multiple items. We also label these four types of tokens in Figure 1. Below, we first describe how to construct the ActionPiece tokenizer’s vocabulary given a corpus of action sequences (Section 3.2.1). Then, we introduce how to segment action sequences into a new sequence of sets, where each set corresponds to a token from the constructed vocabulary (Section 3.2.2).
@@ 3.2 Contextual Action Sequence Tokenizer
7:    Select $(c_{u},c_{v})\leftarrow\operatorname*{arg\,max}_{(c_{i},c_{j})}\text{count}(c_{i},c_{j})$
@@ 3.2.1 Vocabulary Construction on Action Sequence Corpus
Update: corpus updating with action-intermediate nodes. The next step in each iteration is to merge the token pair with the highest accumulated co-occurrence weight. Since token merging may change the set size, we use a double-ended linked list (Zouhar et al., 2023) to maintain the action sequences, where each node represents a set of tokens. Merging tokens within the same set is straightforward, i.e., replacing the two tokens with a new one. However, merging tokens from two adjacent sets is more complex, e.g., determining which set should include the new token.
Intermediate Node. We introduce the concept of “intermediate node” to handle tokens that combine features from multiple sets. Initially, all nodes in the maintained linked list contain features specific to their corresponding actions. These nodes are referred to as “action nodes.”
@@
Data structures. To address this, we propose maintaining a hash table to store the accumulated co-occurrences for each token pair. Additionally, we use inverted indices to map token pairs to all the linked lists that contain them. A global heap is maintained to return the token pair with the highest accumulated co-occurrence. The key challenge lies in updating these data structures. We carefully compute the changes in accumulated co-occurrences and update both the hash table and the inverted indices. For the heap, we employ a lazy-update strategy. We insert the latest weights with a tag. When fetching a value from the heap, we check the tag to verify if the value is up-to-date. If it is not, we discard the value and fetch the next one.
@@
3.2.2 Segmentation by Set Permutation Regularization
3.3 Generative Recommendation Models
3.3.1 Training on Augmented Token Sequences
3.3.2 Inference-Time Ensembling
@@ 4.1 Experimental Setup
Datasets. We use three categories from the Amazon Reviews dataset (McAuley et al., 2015) for our experiments: “Sports and Outdoors” (Sports), “Beauty” (Beauty), and “CDs and Vinyl” (CDs). Each user’s historical reviews are considered “actions” and are sorted chronologically as action sequences, with earlier reviews appearing first. To evaluate the models, we adopt the widely used leave-last-out protocol (Kang & McAuley, 2018; Zhao et al., 2022; Rajput et al., 2023), where the last item and second-to-last item in each action sequence are used for testing and validation, respectively. More details about the datasets can be found in Appendix E.
Compared methods. We compare the performance of ActionPiece with the following methods: (1) ID-based sequential recommendation methods, including BERT4Rec (Sun et al., 2019), and SASRec (Kang & McAuley, 2018); (2) feature-enhanced sequential recommendation methods, such as FDSA (Zhang et al., 2019), S3-Rec (Zhou et al., 2020), and VQ-Rec (Hou et al., 2023); and (3) generative recommendation methods, including P5-CID (Hua et al., 2023), TIGER (Rajput et al., 2023), LMIndexer (Jin et al., 2024), HSTU (Zhai et al., 2024), and SPM-SID (Singh et al., 2024), each representing a different action tokenization method (Table 1). A detailed description of these baselines is provided in Appendix F.
@@
Implementation details. Please refer to Appendix G for detailed implementation and hyperparameter settings.
@@ 4.2 Overall Performance
TIGER with larger vocabularies
@@
not applicable as $2^{14}$ is larger than #items in Beauty.
@@ 4.3 Ablation Study
    (1) We increase the vocabulary size of TIGER, to determine whether the performance gain of ActionPiece is solely due to scaling up the number of tokens in the vocabulary. By increasing the number of semantic ID digits per item ($4\rightarrow 6$) and the number of candidate semantic IDs per digit ($2^{8}\rightarrow 2^{13}\ \text{or}\ 2^{14}$), we create two variants with vocabularies larger than ActionPiece. However, these TIGER variants perform worse than ActionPiece, and even the original TIGER with only $1024$ tokens. The experimental results suggest that scaling up the vocabulary size for generative recommendation models is challenging, consistent with the observations from Zhang et al. (2024).
@@
    (3) To evaluate the effectiveness of SPR, we revert to naive segmentation, as described in Section 3.2.2, during model training and inference, respectively. The results show that replacing SPR with naive segmentation in either training or inference degrades performance.
@@ 5 Conclusion
$\mathcal{V}$ vocabulary of ActionPiece tokenizer
$\mathcal{R}$ merge rules of ActionPiece tokenizer
@@
$Q=|\mathcal{V}|$ size of ActionPiece vocabulary
$P(c,c^{\prime})$ probability that tokens $c$ and $c^{\prime}$ are adjacent when flattening a sequence of sets into a token sequence

### v2

@@ Abstract
Generative recommendation (GR) is an emerging paradigm where user actions are tokenized into discrete token patterns and autoregressively generated as predictions. However, existing GR models tokenize each action independently, assigning the same fixed tokens to identical actions across all sequences without considering contextual relationships. This lack of context-awareness can lead to suboptimal performance, as the same action may hold different meanings depending on its surrounding context. To address this issue, we propose ActionPiece to explicitly incorporate context when tokenizing action sequences. In ActionPiece, each action is represented as a set of item features. Given the action sequence corpora, we construct the vocabulary by merging feature patterns as new tokens, based on their co-occurrence frequency both within individual sets and across adjacent sets. Considering the unordered nature of feature sets, we further introduce set permutation regularization, which produces multiple segmentations of action sequences with the same semantics. Our code is available at: https://github.com/google-deepmind/action_piece.
@@ 1 Introduction
Generative recommendation (GR) (Rajput et al., 2023; Zheng et al., 2024; Zhai et al., 2024) is an emerging paradigm for the sequential recommendation task (Hidasi et al., 2016; Kang & McAuley, 2018). By tokenizing the user actions (typically represented by the interacted items) into discrete tokens, GR models learn to autoregressively generate tokens, which are then parsed into recommended items. These tokens share a compact vocabulary that does not scale with the item pool size, improving model scalability, memory efficiency, and recommendation performance. The input action sequence is vital in understanding user intentions (Hidasi et al., 2016; Li et al., 2017; Kang & McAuley, 2018), which organizes a user’s historical interactions in chronological order. The same action (e.g., purchasing the same item) may have different meanings in different action sequences. Evidence of taking a certain action can be found in the context, such as whether other items in the sequence share the same brand, color tone, or price range (Zhang et al., 2019; Zhou et al., 2020; Hou et al., 2022, 2023; Yuan et al., 2023).
Figure 1: Illustration of the tokenization process of ActionPiece. Each action is represented as an unordered feature set. This figure presents two possible tokenized sequences, where features are grouped into different segments. The same action can be tokenized into different tokens depending on the surrounding context. A detailed case study can be found in Section 4.5.
Despite the importance of contextual relations among actions, existing methods tokenize each action independently of its context (summarized in  Table 1). The typical pipeline for tokenizing action sequences involves two steps: (1) Tokenizing each action/item individually into a pattern of tokens; (2) Replacing each action in the input sequence with its corresponding token pattern. In this way, the tokens do not explicitly contain the context. Instead, they solely rely on the autoregressive model’s parameters being well-trained to generalize effectively in understanding the context, which challenges the capabilities of GR models. As a comparison, tokenization in language modeling also originates from context-independent methods, such as word-level tokenization (Sutskever et al., 2014; Bahdanau et al., 2015). A decade of progress has led to most tokenization methods for modern large language models (LLMs) (OpenAI, 2022; Anil et al., 2023; Touvron et al., 2023; Zhao et al., 2023) adopting context-aware approaches, including BPE (Sennrich et al., 2016) and Unigram tokenization (Kudo, 2018), which tokenize the same word roots along with their adjacent context into different tokens.
@@
To this end, we propose ActionPiece, which enables the same actions to be tokenized into different tokens based on their surrounding context. (1) Vocabulary construction begins by initializing the vocabulary to include every unique feature as initial tokens. The vocabulary is then constructed by iteratively learning merge rules. Each merge rule specifies that a pair of tokens can be merged into a new token. In each iteration, we enumerate the training corpus to count the co-occurrence of existing tokens. Considering the structural differences between token pairs, e.g., whether they occur within a single set or between two adjacent sets, we assign different weights to different pairs. (2) Segmentation refers to dividing raw features in action sequences into groups that can be replaced by tokens from the vocabulary. To fully exploit the unordered nature of the feature set of each action, we introduce set permutation regularization. By randomly permuting the features within each set, we can produce multiple token sequences of a single action sequence that preserve the same semantics. These variations act as natural augmentations for training data and enable inherent ensembling during model inference.
@@ 2 Related Work
Generative recommendation. Conventional sequential recommendation models often relies on large embedding tables to store representations for all items, leading to significant engineering and optimization challenges (Hidasi et al., 2016; Li et al., 2017; Kang & McAuley, 2018). Generative recommendation (Rajput et al., 2023; Zheng et al., 2024; Zhai et al., 2024; Deldjoo et al., 2024; Hou et al., 2025) addresses these issues by tokenizing each item as tokens from a shared vocabulary. By autoregressively generating the next tokens as recommendations, this generative paradigm offers benefits such as memory efficiency (Rajput et al., 2023; Yang et al., 2024; Ding et al., 2024), scalability (Zhai et al., 2024; Liu et al., 2024b), and easier alignment with LLMs (Zheng et al., 2024; Jin et al., 2024; Tan et al., 2024; Li et al., 2025). Existing research has developed different action tokenization techniques, such as hierarchical clustering (Hua et al., 2023; Si et al., 2024), quantization (Rajput et al., 2023; Wang et al., 2024a; Zhu et al., 2024a), or jointly training with recommendation models (Liu et al., 2025). Other works incorporate additional modalities like collaborative filtering (Petrov & Macdonald, 2023; Wang et al., 2024c, b; Liu et al., 2024b, a) and natural language (Zheng et al., 2024; Jin et al., 2024; Hou et al., 2024b; Zhang et al., 2025a). However, current methods tokenize each action independently, ignoring the surrounding context. In this work, we propose the first context-aware action tokenization method, where the same actions are tokenized differently in different action sequences.
@@
7:    Select $(c_{u},c_{v})\leftarrow\arg\max_{(c_{i},c_{j})}\text{count}(c_{i},c_{j})$
@@
3 Method
In this section, we present ActionPiece, a context-aware method for tokenizing action sequences for generative recommendation. First, we formulate the task in Section 3.1. Then, we introduce the proposed tokenizer, covering vocabulary construction and segmentation, in Section 3.2. Finally, we describe the model training and inference process using ActionPiece-tokenized sequences in Section 3.3.
3.1 Problem Formulation
3.2 Contextual Action Sequence Tokenizer
The proposed tokenizer is designed to transform action sequences (represented as sequences of feature sets) into token sequences. In the ActionPiece-tokenized sequences, each token corresponds to a set containing varying numbers of features. For example, a token can represent: (1) a subset of features from one item; (2) a single feature; (3) all features of one item; or (4) features from multiple items. We also label these four types of tokens in Figure 1. Below, we first describe how to construct the ActionPiece tokenizer’s vocabulary given a corpus of action sequences (Section 3.2.1). Then, we introduce how to segment action sequences into a new sequence of sets, where each set corresponds to a token from the constructed vocabulary (Section 3.2.2).
@@ 3.2.1 Vocabulary Construction on Action Sequence Corpus
Update: corpus updating with action-intermediate nodes. The next step in each iteration is to merge the token pair with the highest accumulated co-occurrence weight. Since token merging may change the set size, we use a double-ended linked list (Zouhar et al., 2023) to maintain each action sequence, where each node represents a set of tokens. Merging tokens within the same set is straightforward, i.e., replacing the two tokens with a new one. However, merging tokens from two adjacent sets is more complex, e.g., determining which set should include the new token.
Intermediate Node. We introduce the concept of “intermediate node” to handle tokens that combine features from multiple sets. Initially, all nodes in the maintained linked lists contain features specific to their corresponding actions. These nodes are referred to as “action nodes.”
@@
Data structures. To address this, we propose creating inverted indices to map token pairs to all the linked lists that contain them. A global heap is maintained to return the token pair with the highest accumulated co-occurrence. The key challenge lies in updating these data structures. We carefully compute the changes in accumulated co-occurrences and update the inverted indices. For the heap, we employ a lazy-update strategy. We insert the latest weights with a tag. When fetching a value from the heap, we check the tag to verify if the value is up-to-date. If it is not, we discard the value and fetch the next one.
@@
3.2.2 Segmentation by Set Permutation Regularization
Table 2: Comparison between the widely used text tokenization method, byte-pair encoding (BPE) (Sennrich et al., 2016), and the proposed context-aware action sequence tokenization method, ActionPiece.
Aspect BPE ActionPiece
Data Type text sequences action (unordered feature set) sequences
Token a byte sequence a feature set
Initial Vocabulary single bytes single-feature sets
3.3 Generative Recommendation Models
3.3.1 Training on Augmented Token Sequences
3.3.2 Inference-Time Ensembling
3.4 Discussion
Orders in action sequences. In recommender systems, user actions are typically represented as sets of features, such as the title and price of the associated item. These features typically have no inherent order within a single action. However, in sequential recommendation tasks, a user’s historical actions are usually ordered by timestamp to capture temporal behavioral dynamics. Building on this, we model action sequences as sequences of feature sets: while the features within each action remain unordered, the temporal ordering of actions is preserved. This evolving composition of features over time captures meaningful sequential patterns.
ActionPiece vs. BPE. While ActionPiece follows a similar algorithmic framework as BPE, the key distinction lies in the data formats they are designed to model. BPE operates on one-dimensional byte sequences, whereas ActionPiece is tailored for tokenizing sequences of feature sets. Modeling each action as an unordered set aligns better with the inherent structure of action sequences. For clarity, we summarize the key differences in Table 2.
Efficiency impact of SPR. Despite introducing set permutation regularization, the training efficiency remains comparable to existing methods such as TIGER. Feature permutation is performed on the CPU and runs asynchronously alongside TPU/GPU-based model updates, resulting in no noticeable degradation in training speed. At inference time, SPR introduces additional FLOPs due to the ensemble of augmented test cases. However, the overall latency remains comparable to baseline methods, as the augmented versions can be processed in parallel across multiple computing devices (e.g., TPUs or GPUs). This parallelism offsets the added computation, enabling our method to maintain efficient inference despite the use of inference-time ensembling.
Datasets #Users #Items #Actions Avg. $t$
Sports 18,357 35,598 260,739 8.32
Beauty 22,363 12,101 176,139 8.87
CDs 75,258 64,443 1,022,334 14.58
@@ 4.1 Experimental Setup
Datasets. We use three categories from the Amazon Reviews dataset (McAuley et al., 2015) for our experiments: “Sports and Outdoors” (Sports), “Beauty” (Beauty), and “CDs and Vinyl” (CDs). Each user’s historical reviews are considered “actions” and are sorted chronologically as action sequences, with earlier reviews appearing first. To evaluate the models, we adopt the widely used leave-last-out protocol (Kang & McAuley, 2018; Zhao et al., 2022; Rajput et al., 2023), where the last item and second-to-last item in each action sequence are used for testing and validation, respectively. The statistics of the processed datasets are shown in Table 3. More details about the datasets can be found in Appendix F.
Compared methods. We compare the performance of ActionPiece with the following methods: (1) ID-based sequential recommendation methods, including BERT4Rec (Sun et al., 2019), and SASRec (Kang & McAuley, 2018); (2) feature-enhanced sequential recommendation methods, such as FDSA (Zhang et al., 2019), S3-Rec (Zhou et al., 2020), and VQ-Rec (Hou et al., 2023); and (3) generative recommendation methods, including P5-CID (Hua et al., 2023), TIGER (Rajput et al., 2023), LMIndexer (Jin et al., 2024), HSTU (Zhai et al., 2024), and SPM-SID (Singh et al., 2024), each representing a different action tokenization method (Table 1). A detailed description of these baselines is provided in Appendix G.
@@
Implementation details. Please refer to Appendix H for detailed implementation and hyperparameter settings.
@@ 4.2 Overall Performance
TIGER with varying vocabulary sizes
(1.1) TIGER - 192 ($4\times 48$) 0.0231 0.0362 N/A
(1.2) TIGER - 768 ($3\times 2^{8}$) 0.0220 0.0378 0.0331
@@
not applicable because the number of conflicts among semantic ID prefixes (the first three tokens) in CDs exceeds 48.
not applicable because $2^{14}$ is larger than #items in Beauty.
@@ 4.3 Ablation Study
    (1) To examine whether the performance gain of ActionPiece stems from the choice of vocabulary size, we conduct an ablation study by varying the vocabulary size of TIGER. We increase the number of semantic ID digits per item ($4\rightarrow 6$) and the number of candidate semantic IDs per digit ($2^{8}\rightarrow 2^{13}$ or $2^{14}$), resulting in two TIGER variants with vocabularies larger than ActionPiece. We also include two variants with reduced vocabulary sizes by decreasing the number of digits or candidates per digit. Despite the broader range of vocabulary sizes, all TIGER variants perform worse than ActionPiece, and in some cases, even worse than the original TIGER with 1,024 tokens. These results suggest that the performance improvement of ActionPiece is not simply due to scaling the vocabulary size up or down. Instead, they highlight the difficulty of effectively scaling vocabulary size in generative recommendation models, consistent with the observations from Zhang et al. (2024).
@@
    (3) To evaluate the effectiveness of SPR, we revert to naive segmentation, as described in Section 3.2.2, during model training and inference, respectively. The results show that replacing SPR with naive segmentation in either training or inference degrades performance. We also introduce an ablation variant that applies SPR to the existing GR model TIGER. The results indicate that SPR alone is insufficient to improve the GR model. When applied to action sequences tokenized by context-independent methods, SPR does not change token frequencies and merely disrupts the internal token order within RQ-VAE-based semantic IDs. In contrast, ActionPiece derives different tokens for the same item based on its neighboring context, improving token utilization and serving as an effective form of data augmentation.
@@ 5 Conclusion
Acknowledgements
This work was conducted while Yupeng Hou was a student researcher at Google DeepMind. We thank Yichen Zhou for helpful discussions and Nikhil Mehta for valuable suggestions on the draft.
Impact Statement
This paper introduces ActionPiece, a context-aware action sequence tokenizer to enhance generative recommendation. This work aims to advance personalized recommender systems, enabling more accurate understanding of user preferences. We argue that this work is not directly correlated to certain society or ethical concerns. The impact of this work is primarily tied to the broader implications of recommender systems in various domains, such as e-commerce, entertainment, and social platforms.
@@ Acknowledgements
$\mathcal{V}$ The vocabulary of ActionPiece tokenizer
$\mathcal{R}$ The merge rules of ActionPiece tokenizer
@@
$Q=|\mathcal{V}|$ The size of ActionPiece vocabulary
$P(c,c^{\prime})$ The probability that tokens $c$ and $c^{\prime}$ are adjacent when flattening a sequence of sets into a token sequence